# Variable Selection in Convex Piecewise Linear Regression

### v1

@@ 5 Discussion
$\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\theta}^{t+1}_{j}-\bm{\alpha}_{j}^{t+1}\right)\right\|_{2}\leq\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{\alpha}_{j}^{t+1}-\bm{\theta}_{j}^{\star}\right)\right\|_{2},\quad\forall j\in[k]$
which follows from the fact that $\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\theta}_{j}^{t+1}=\bm{\theta}_{j}^{t+1}$ coincides with the projection of $\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\alpha}_{j}^{t+1}$ onto $\Gamma_{s}$, and $\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\theta}_{j}^{\star}=\bm{\theta}_{j}^{\star}$ belongs to $\Gamma_{s}$ for all $j\in[k]$.
@@
$\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\varphi^{t}}\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\varphi^{t}\cap\varphi^{t+1}}\bm{h}_{j}^{t}+\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}_{j}^{t}+\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}_{j}^{t}.$ (42)
@@
$\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\left(\langle\bm{\xi}_{i},\bm{\theta}_{j}^{t}\rangle-\max_{j\in[k]}\langle\bm{\xi}_{i},\bm{\theta}_{j}^{\star}\rangle-z_{i}\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}.$ (43)
@@
$=\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\xi}_{i},\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}_{j}^{t}\rangle+\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\xi}_{i},\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{h}_{j}^{t}\rangle+\sum_{j^{\prime}\neq j}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}^{\star}_{j^{\prime}}\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle$ (44)
@@
$\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})$ $=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i},\bm{h}^{t}_{j}\rangle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{p}_{j}}\quad+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\widetilde{\Pi}}_{\varphi^{t}\backslash\varphi^{t+1}}\bm{\xi}_{i},\bm{h}^{t}_{j}\rangle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{q}_{j}}$ (45)
$\quad+\underbrace{\frac{1}{n}\sum_{\begin{subarray}{c}i=1\\ j^{\prime}\neq j\end{subarray}}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{c}_{j}}\quad-\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}z_{i}\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\xi}_{i}}_{\bm{d}_{j}}.$
@@
$\frac{1}{4}\left\|\bm{h}^{t+1}\right\|_{2}^{2}$ $\leq\sum_{j=1}^{k}\left[\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{h}^{t}_{j}-\mu_{j}^{t}\bm{p}_{j}\right)\right\|_{2}+\mu_{j}^{t}\left(\left\|\bm{q}_{j}\right\|_{2}+\left\|\bm{c}_{j}\right\|_{2}+\left\|\bm{d}_{j}\right\|_{2}\right)\right]^{2}.$ (46)
@@
Instate the assumptions and definitions in Theorem 2.3. Then, the following events hold jointly for all $t\in\mathbb{N}\cup\{0\}$ and $j\in[k]$ with probability at least $1-\delta$:
$\sup\limits_{\begin{subarray}{c}|\varphi|\leq s\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\varphi}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi}\right\|\leq k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})},$ (47)
$\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}-\pi_{j}^{t}\right|\leq k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})},$ (48)
@@
$\sup\limits_{\begin{subarray}{c}|\varphi|\leq s\end{subarray}}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}z_{i}\bm{\widetilde{\Pi}}_{\varphi}[\bm{x}_{i};1]\right\rVert}_{2}\leq C\sigma_{z}\sqrt{\frac{sk\log\left({n}/{s}\right)+s\log\left({d}/{s}\right)+\log\left({1}/{\delta}\right)}{n}}\triangleq\tau_{\mathrm{noise}},$ (50)
$\sup\limits_{\begin{subarray}{c}|\varphi|\leq s\end{subarray}}{\left\lVert\frac{1}{\sqrt{n}}\bm{\widetilde{\Pi}}_{\varphi}[\bm{\xi}_{1},\ldots,\bm{\xi}_{n}]\right\rVert}\leq k^{-2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$ (51)
where $C>0$ is some absolute constant.
We proceed with the proof under the assumption that all the statements in Lemma A.1 hold. For notational simplicity, we will use the short hand $\epsilon_{\mathrm{min}}\triangleq k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$ repeatedly. We can now begin upper bounding the terms in (46). The first summand in (46) is upper-bounded as
$\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{h}^{t}_{j}-\mu\bm{p}_{j}\right)\|_{2}$
$\leq{\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{I}_{d+1}-\frac{\mu_{j}k}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}\}}\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right\rVert}\cdot{\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\right\rVert}_{2}.$ (52)
@@
${\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{\left missing}(\bm{I}_{d+1}-\frac{\mu^{t}_{j}}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right\rVert}$
$\leq\mu^{t}_{j}\left(\underbrace{\left\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\right\|}_{\mathcal{A}}+\underbrace{\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}-\frac{1}{\mu^{t}_{j}}\right|}_{\mathcal{B}}\right).$ (53)
By (47) we have that $\mathcal{A}\leq\epsilon_{\mathrm{min}}$, $\mathcal{B}=0$ by the definition of $\mu_{j}$. Also, since $\bm{\theta}^{t}\in\mathcal{N}(\bm{\theta}^{\star},\sqrt{2}\Delta\rho)$, by Lemma B.3 we have that
@@
$\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\left(\bm{h}^{t}_{j}-\mu_{j}^{t}\bm{p}_{j}\right)\|_{2}\leq\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}{\left\lVert\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\right\rVert}_{2}.$ (56)
We can use (47) again to bound the second summand of (46) as
$\mu^{t}_{j}\|\bm{q}_{j}\|_{2}\leq\mu^{t}_{j}\epsilon\|\bm{\widetilde{\Pi}}_{\varphi^{t+1}}\bm{h}^{t}_{j}\|_{2}\leq\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}.$
The vector $\bm{c}_{j}$ in the last term of (45) is factorized as $\bm{c}_{j}=\bm{\widetilde{\Pi}}_{\varphi^{t+1}}[\bm{\xi}_{1},\ldots,\bm{\xi}_{n}]\bm{v}$, where
$\bm{v}\triangleq\sum_{j^{\prime}\neq j}\begin{bmatrix}\mathbb{1}_{\{\bm{x}_{1}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{1},\bm{v}_{jj^{\prime}}^{\star}\rangle\\ \vdots\\ \mathbb{1}_{\{\bm{x}_{n}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{n},\bm{v}_{jj^{\prime}}^{\star}\rangle\end{bmatrix}.$
Therefore, we have
$\|\bm{c}_{j}\|_{2}\leq{\left\lVert\frac{1}{\sqrt{n}}\bm{\widetilde{\Pi}}_{\varphi^{t+1}}[\bm{\xi}_{1},\ldots,\bm{\xi}_{n}]\right\rVert}\cdot{\left\lVert\frac{1}{\sqrt{n}}\bm{v}\right\rVert}_{2}\leq\frac{\epsilon_{\mathrm{min}}}{\sqrt{k}}{\left\lVert\frac{1}{\sqrt{n}}\bm{v}\right\rVert}_{2},$ (57)
@@
$\leq\underbrace{\frac{2}{5\gamma}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}k^{-1}}_{\lambda^{2}}\sum_{j^{\prime}:j^{\prime}\neq j}\|\bm{v}_{jj^{\prime}}^{t}-\bm{v}_{jj^{\prime}}^{\star}\|_{2}^{2}$
$=\lambda^{2}\sum_{j^{\prime}:j^{\prime}\neq j}\|\bm{h}_{j}^{t}-\bm{h}_{j^{\prime}}^{t}\|_{2}^{2}$
$\leq\lambda^{2}\sum_{j^{\prime}:j^{\prime}\neq j}\left(\|\bm{h}_{j}^{t}\|_{2}^{2}+\|\bm{h}_{j^{\prime}}^{t}\|_{2}^{2}\right),$ (58)
where the first inequality follows from (49). Since the $\ell_{1}$ norm dominates the $\ell_{2}$ norm, we can write (A) as
$\|\frac{1}{\sqrt{n}}\bm{v}\|_{2}\leq\lambda\sum_{j^{\prime}:j^{\prime}\neq j}\left(\|\bm{h}_{j}^{t}\|_{2}+\|\bm{h}_{j^{\prime}}^{t}\|_{2}\right).$ (59)

### v2

@@ 6 Discussion
$\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\theta}^{t+1}_{j}-\bm{\alpha}_{j}^{t+1}\right)\right\|_{2}\leq\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\alpha}_{j}^{t+1}-\bm{\theta}_{j}^{\star}\right)\right\|_{2},\quad\forall j\in[k]$
which follows from the fact that $\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\theta}_{j}^{t+1}=\bm{\theta}_{j}^{t+1}$ coincides with the projection of $\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\alpha}_{j}^{t+1}$ onto $\Gamma_{s}$, and $\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\theta}_{j}^{\star}=\bm{\theta}_{j}^{\star}$ belongs to $\Gamma_{s}$ for all $j\in[k]$.
@@
$\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}}\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}\cap\mathcal{U}^{t+1}}\bm{h}_{j}^{t}+\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}\backslash\mathcal{U}^{t+1}}\bm{h}_{j}^{t}=\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{h}_{j}^{t}+\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}\backslash\mathcal{U}^{t+1}}\bm{h}_{j}^{t}.$ (49)
@@
$\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\left(\langle\bm{\xi}_{i},\bm{\theta}_{j}^{t}\rangle-\max_{j\in[k]}\langle\bm{\xi}_{i},\bm{\theta}_{j}^{\star}\rangle-z_{i}\right)\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\xi}_{i}.$ (50)
@@
$=\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\xi}_{i},\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{h}_{j}^{t}\rangle+\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\xi}_{i},\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}\backslash\mathcal{U}^{t+1}}\bm{h}_{j}^{t}\rangle+\sum_{j^{\prime}\neq j}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}^{\star}_{j^{\prime}}\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle$ (51)
@@
$\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\nabla_{\bm{\theta}_{j}}\ell(\bm{\theta}^{t})$ $=\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\xi}_{i},\bm{h}^{t}_{j}\rangle\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\xi}_{i}}_{\bm{p}_{j}}$ (52)
$\quad+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{U}^{t}\backslash\mathcal{U}^{t+1}}\bm{\xi}_{i},\bm{h}^{t}_{j}\rangle\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\xi}_{i}}_{\bm{q}_{j}}$
$\quad+\underbrace{\frac{1}{n}\sum_{\begin{subarray}{c}i=1\\ j^{\prime}\neq j\end{subarray}}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\xi}_{i},\bm{v}_{jj^{\prime}}^{\star}\rangle\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\xi}_{i}}_{\bm{c}_{j}}\quad-\underbrace{\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}z_{i}\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\xi}_{i}}_{\bm{d}_{j}}.$
@@
$\frac{1}{4}\left\|\bm{h}^{t+1}\right\|_{2}^{2}$ $\leq\sum_{j=1}^{k}\left[\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{h}^{t}_{j}-\mu_{j}^{t}\bm{p}_{j}\right)\right\|_{2}+\mu_{j}^{t}\left(\left\|\bm{q}_{j}\right\|_{2}+\left\|\bm{c}_{j}\right\|_{2}+\left\|\bm{d}_{j}\right\|_{2}\right)\right]^{2}.$ (53)
@@
Instate the assumptions and definitions in Theorem 2.3, and let $\epsilon_{\mathrm{min}}\triangleq k^{-3/2}\pi_{\mathrm{min}}^{2(1+\zeta^{-1})}$. Then, the following events hold jointly for all $t\in\mathbb{N}\cup\{0\}$ and $j\in[k]$ with probability at least $1-\delta$:
$\sup\limits_{\begin{subarray}{c}|\mathcal{U}|\leq s\end{subarray}}\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\mathcal{U}}\right\|\leq\epsilon_{\mathrm{min}},$ (54)
$\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}-\pi_{j}^{t}\right|\leq\epsilon_{\mathrm{min}},$ (55)
@@
$\sup\limits_{\begin{subarray}{c}|\mathcal{U}|\leq s\end{subarray}}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}^{t}\}}z_{i}\bm{\widetilde{\Pi}}_{\mathcal{U}}[\bm{x}_{i};1]\right\rVert}_{2}\leq C\sigma_{z}\sqrt{\frac{sk\log\left({n}/{s}\right)+s\log\left({d}/{s}\right)+\log\left({1}/{\delta}\right)}{n}}\triangleq\tau_{\mathrm{noise}},$ (57)
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sup\limits_{\begin{subarray}{c}|\mathcal{U}|\leq s\end{subarray}}{\left\lVert\frac{1}{\sqrt{n}}\bm{\widetilde{\Pi}}_{\mathcal{U}}[\mathbb{1}_{\{\bm{x}_{1}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\bm{\xi}_{1},\ldots,\mathbb{1}_{\{\bm{x}_{n}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\bm{\xi}_{n}]\right\rVert}\leq{\frac{\pi_{\min}^{(1+\zeta^{-1})/2}}{k^{1/2}}}\quad\forall j^{\prime}\neq j.$} (58)
The proof of all statements in this lemma is deferred to Appendix B. We proceed with the proof under the assumption that all the statements in Lemma A.1 hold. We can now begin upper bounding the terms in (53). The first summand in (53) is upper-bounded as
$\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{h}^{t}_{j}-\mu\bm{p}_{j}\right)\|_{2}\leq{\left\lVert\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{I}_{d+1}-\frac{\mu_{j}k}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}_{j}\}}\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}\right)\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\right\rVert}\cdot{\left\lVert\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{h}^{t}_{j}\right\rVert}_{2}.$ (59)
@@
${\left\lVert\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{\left.}(\bm{I}_{d+1}-\frac{\mu^{t}_{j}}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}\right)\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\right\rVert}$
$\leq\mu^{t}_{j}\left(\underbrace{\left\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\scriptscriptstyle{\textup{{T}}}}-\bm{I}_{d+1}\right)\right)\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\right\|}_{\mathcal{A}}+\underbrace{\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}_{i}\in\mathcal{C}^{t}_{j}\}}-\frac{1}{\mu^{t}_{j}}\right|}_{\mathcal{B}}\right).$ (60)
By (54) we have that $\mathcal{A}\leq\epsilon_{\mathrm{min}}$. Furthermore, $\mathcal{B}=0$ by the definition of $\mu_{j}$. Also, since $\bm{\theta}^{t}\in\mathcal{N}(\bm{\theta}^{\star},\sqrt{2}\Delta\rho)$, by Lemma B.4 we have that
@@
$\|\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{h}^{t}_{j}-\mu_{j}^{t}\bm{p}_{j}\right)\|_{2}\leq\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}{\left\lVert\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}\bm{h}^{t}_{j}\right\rVert}_{2}.$ (63)
The second summand of (53) is written as
$\mu^{t}_{j}\|\bm{q}_{j}\|_{2}$ $=\mu^{t}_{j}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}\in\mathcal{C}_{j}^{t}\}}\widetilde{\bm{\Pi}}_{\mathcal{U}^{t+1}}\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}\widetilde{\bm{\Pi}}_{\mathcal{U}^{t}\setminus\mathcal{U}^{t+1}}\bm{h}_{j}^{t}\right\rVert}$
$=\mu^{t}_{j}{\left\lVert\left[\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}\in\mathcal{C}_{j}^{t}\}}\widetilde{\bm{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}-\bm{I}_{d+1}\right)\widetilde{\bm{\Pi}}_{\mathcal{U}^{t}\setminus\mathcal{U}^{t+1}}\right]\widetilde{\bm{\Pi}}_{\mathcal{U}^{t}\setminus\mathcal{U}^{t+1}}\bm{h}_{j}^{t}\right\rVert}$
$\leq\frac{1}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}{\left\lVert\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}_{\{\bm{x}\in\mathcal{C}_{j}^{t}\}}\widetilde{\bm{\Pi}}_{\mathcal{U}^{t+1}}\left(\bm{\xi}_{i}\bm{\xi}_{i}^{\mathsf{T}}-\bm{I}_{d+1}\right)\widetilde{\bm{\Pi}}_{\mathcal{U}^{t}\setminus\mathcal{U}^{t+1}}\right\rVert}\cdot{\left\lVert\widetilde{\bm{\Pi}}_{\mathcal{U}^{t}\setminus\mathcal{U}^{t+1}}\bm{h}_{j}^{t}\right\rVert}$
$\leq\frac{\epsilon_{\mathrm{min}}}{(1-\varrho)\pi^{\star}_{j}-\epsilon_{\mathrm{min}}}{\left\lVert\widetilde{\bm{\Pi}}_{\mathcal{U}^{t}\setminus\mathcal{U}^{t+1}}\bm{h}_{j}^{t}\right\rVert},$ (64)
where the second equality follows from the idempotency of projection matrices and the observation that $\widetilde{\bm{\Pi}}_{\mathcal{U}^{t+1}}\bm{I}_{d+1}\widetilde{\bm{\Pi}}_{\mathcal{U}^{t}\setminus\mathcal{U}^{t+1}}=\bm{0}$, the first inequality follows from the upper bound on $\mu_{j}^{t}$ in (62) and the definition of the operator norm, and the last inequality follows from (54).
The vector $\bm{c}_{j}$ in the last term of (52) is factorized as $\bm{c}_{j}=\frac{1}{n}\bm{E}\bm{v}$, where
$\bm{v}\triangleq\sum_{j^{\prime}\neq j}\begin{bmatrix}\mathbb{1}_{\{\bm{x}_{1}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{1},\bm{v}_{jj^{\prime}}^{\star}\rangle\\ \vdots\\ \mathbb{1}_{\{\bm{x}_{n}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\langle\bm{\widetilde{\Pi}}_{\mathcal{S}^{\star}}\bm{\xi}_{n},\bm{v}_{jj^{\prime}}^{\star}\rangle\end{bmatrix}$,
and $\bm{E}=\bm{\widetilde{\Pi}}_{\mathcal{U}^{t+1}}[\mathbb{1}_{\{\bm{x}_{1}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\bm{\xi}_{1},\ldots,\mathbb{1}_{\{\bm{x}_{n}\in\mathcal{C}^{t}_{j}\cap\mathcal{C}_{j^{\prime}}^{\star}\}}\bm{\xi}_{n}]$. Therefore, we have
$\|\bm{c}_{j}\|_{2}\leq{\left\lVert\frac{1}{\sqrt{n}}\bm{E}\right\rVert}\cdot{\left\lVert\frac{1}{\sqrt{n}}\bm{v}\right\rVert}_{2}\leq{\frac{\pi_{\min}^{(1+\zeta^{-1})/2}}{k^{1/2}}}{\left\lVert\frac{1}{\sqrt{n}}\bm{v}\right\rVert}_{2},$ (65)
@@
$\leq\frac{2}{5\gamma}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}k^{-1}\sum_{j^{\prime}:j^{\prime}\neq j}\|\bm{v}_{jj^{\prime}}^{t}-\bm{v}_{jj^{\prime}}^{\star}\|_{2}^{2}$
$=\frac{2}{5\gamma}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}k^{-1}\sum_{j^{\prime}:j^{\prime}\neq j}\|\bm{h}_{j}^{t}-\bm{h}_{j^{\prime}}^{t}\|_{2}^{2}$
$\leq\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4}}{5\gamma}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}k^{-1}\sum_{j^{\prime}:j^{\prime}\neq j}\left(\|\bm{h}_{j}^{t}\|_{2}^{2}+\|\bm{h}_{j^{\prime}}^{t}\|_{2}^{2}\right),$ (66)
where the first equality follows from the non-overlapping property of set partitions, i.e. $\mathcal{C}_{j}^{t}\cap\mathcal{C}_{j^{\prime}}^{t}=\mathcal{C}_{q}^{\star}\cap\mathcal{C}_{q^{\prime}}^{\star}=\varnothing$ when $j\neq j^{\prime}$ and $q\neq q^{\prime}$, and the first inequality follows from (56). Since the $\ell_{1}$ norm dominates the $\ell_{2}$ norm, we can write (A) as
${\left\lVert\frac{1}{\sqrt{n}}\bm{v}\right\rVert}_{2}\leq\sqrt{\frac{4}{5\gamma}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}k^{-1}}\sum_{j^{\prime}:j^{\prime}\neq j}\left(\|\bm{h}_{j}^{t}\|_{2}+\|\bm{h}_{j^{\prime}}^{t}\|_{2}\right)\leq\sqrt{\frac{4}{5\gamma}\left(\frac{\pi_{\min}}{16}\right)^{1+\zeta^{-1}}k}\sum_{j^{\prime}=1}^{k}\|\bm{h}_{j^{\prime}}^{t}\|_{2}.$ (67)
Therefore, we have that
${\left\lVert\bm{c}_{j}\right\rVert}_{2}\leq\underbrace{\sqrt{\frac{4\cdot 16^{-(1+\zeta^{-1})}\pi_{\min}^{2(1+\zeta^{-1})}}{5\gamma}}}_{\lambda}\sum_{j^{\prime}=1}^{k}\|\bm{h}_{j^{\prime}}^{t}\|_{2}$ (68)